OpenStudy (anonymous):

The area of a trapezoid is given by the formula A =1/2 (b 1 + b 2)h, where base b 1 is parallel to base b 2 and h is the height. Solve the formula for b 2. Show your work.

OpenStudy (anonymous):

A= 1/2 x (b1 + b2) x h 2Ah = b1 + b2 2Ah -b1 = b2

OpenStudy (anonymous):

thats the answer

OpenStudy (anonymous):

yup b2= 2Ah - b1

OpenStudy (anonymous):

why is it 2AH, isn't it (2A)/h ?

OpenStudy (anonymous):

oh yah...my mistake...

OpenStudy (anonymous):

The correct formula would be: \[\frac{ 2A }{ h } - b1 = b2\]

OpenStudy (anonymous):

thanks